How much work is done when a 40.0kg mass is lifted vertically against gravity to a height of 5.0 meters?
Answer by mananth(16949)

W= F* S
F= force
g= 9.8 m/sec^2
m= 40 kg
..
F= mg
= 40*9.8 N
=39.2 N
..
Distance moved = 5 meters
work done = F * displacement
work done =39.2 * 5 J
=195 J
Work done = 195 J